Show that f:x mapping x/x+1 is one to one mapping


1
Expert's answer
2020-07-04T09:38:10-0400

f(x) = x/(x+1)

For one to one mapping

we prove f(x1)=f(x2) "\to" x1=x2 .

Let

f(x1) = x1/(x1+1)

f(x2)= x2/(x2+1)

now

f(x1)=f(x2)

x1/(x1+1) = x2/(x2+1)

x1(x2+1) = x2(x1+1)

x1x2+x1 = x2x1+x2

x1 = x2

hence it was proved that this function is one to one mapping.